# The Non-Gaussian to Gaussian Transition: Pointwise Heat Kernel Estimates and Optimal Convergence Rates

### 15 Mar 2026 v1

@@ 3.2. $\alpha$-stability estimates of infinitesimal generators
$\left|\delta_{\vartheta\sigma_{1}z}\nabla^{2}p^{(\alpha),\sigma}_{s,t}(y)-\delta_{\vartheta\sigma_{2}z}\nabla^{2}p^{(\alpha),\sigma}_{s,t}(y)\right|\lesssim(1\wedge\|\sigma_{1}-\sigma_{2}\|)(t-s)^{-1/\alpha}|z|\varrho^{(2)}_{\alpha}(t-s,y).$ (3.19)
Similarly, one sees that, when $|z|\leq\varepsilon$,
$\begin{split}&\quad\left|\left[\delta_{\vartheta\sigma_{1}z}\nabla^{2}p^{(\alpha),\sigma}_{s,t}(y)\right]\sigma_{1}z\cdot\sigma_{1}z-\left[\delta_{\vartheta\sigma_{1}z}\nabla^{2}p^{(\alpha),\sigma}_{s,t}(y)\right]\sigma_{2}z\cdot\sigma_{2}z\right|\\ &\lesssim(1\wedge\|\sigma_{1}-\sigma_{2}\|)(t-s)^{-1/\alpha}|z|\varrho^{(2)}_{\alpha}(t-s,y).\end{split}$ (3.20)
(Step 2) Consequently, using (3.19) and (3.20), under the condition (3.1), by (3.7) and (3.8), we have
$|(\widetilde{\mathscr{D}}^{(\alpha),\sigma_{1}}$ $-\widetilde{\mathscr{D}}^{(\alpha),\sigma_{2}})p^{(\alpha),\sigma}(s,t,y)|\lesssim\xi_{s,t}^{(\alpha)}\times(1\wedge\|\sigma_{1}-\sigma_{2}\|)\varrho^{(2)}_{\alpha}(t-s,y)$
$+$ $(1\wedge\|\sigma_{1}-\sigma_{2}\|)\varrho^{(2)}_{\alpha}(t-s,y)\int_{|z|\leq\varepsilon}|t-s|^{-1/\alpha}|z|^{3}\nu^{(\alpha)}({\mathord{{\mathrm{d}}}}z)$
$+$ $\sum_{i=1,2}\int_{|z|>\varepsilon}{\Big(}(|t-s|^{-1/\alpha}|(\sigma_{1}-\sigma_{2})z|)\wedge 1{\Big)}\varrho_{\alpha}(t-s,y+\sigma_{i}z)\nu^{(\alpha)}({\mathord{{\mathrm{d}}}}z)$
Next, we estimate these three terms in turn.
$\bullet$ By (2.19) and (3.18), one sees that
$\Lambda^{(1)}_{\alpha}(s,t,y)\lesssim\xi_{s,t}^{(\alpha)}\times(1\wedge\|\sigma_{1}-\sigma_{2}\|)|t-s|^{-\frac{2}{\alpha}}\varrho_{\alpha}(t-s,y).$
$\bullet$ By (2.13), we get that
$\Lambda_{\alpha}^{(2)}(s,t,y)$ $\leq(1\wedge\|\sigma_{1}-\sigma_{2}\|)|t-s|^{-3/\alpha}\varrho_{\alpha}(t-s,y)\left(\int_{|z|\leq\varepsilon}|z|^{3}\nu^{(\alpha)}({\mathord{{\mathrm{d}}}}z)\right)$
[…]
$\bullet$ Using the condition (3.1), by Lemma 2.3 and (2.14), we obtain that
$\Lambda^{(3)}_{\alpha}(s,t,y$ $)\leq\sum_{i=1,2}\left(\int_{|z|>|t-s|^{1/\alpha}\vee\frac{|y|}{2\kappa_{1}}}+\int_{\frac{|y|}{2\kappa_{1}}\geq|z|>|t-s|^{1/\alpha}}\right)\varrho_{\alpha}(t-s,y+\sigma_{i}z)\nu^{(\alpha)}({\mathord{{\mathrm{d}}}}z)$
[…]
$\,\,\lesssim(2-\alpha)|t-s|^{-1}\varrho_{\alpha}(t-s,y),$
and, similarly,
$\Lambda^{(3)}_{\alpha}(s,t,y)$ $\lesssim\|\sigma_{1}-\sigma_{2}\|\sum_{i=1,2}\int_{|z|>|t-s|^{1/\alpha}\vee\frac{|y|}{2\kappa_{1}}}(t-s)^{-\frac{1}{\alpha}}|z|\varrho_{\alpha}(t-s,y+\sigma_{i}z)\nu^{(\alpha)}({\mathord{{\mathrm{d}}}}z)$
$\qquad+\|\sigma_{1}-\sigma_{2}\|\varrho_{\alpha}(t-s,y)\int_{\frac{|y|}{2\kappa_{1}}\geq|z|>|t-s|^{1/\alpha}}(t-s)^{-\frac{1}{\alpha}}|z|\nu^{(\alpha)}({\mathord{{\mathrm{d}}}}z)$
$\lesssim(2-\alpha)\|\sigma_{1}-\sigma_{2}\|(t-s)^{-1}\left(1+(t-s)^{-1/\alpha}|y|\right)\varrho_{\alpha}(t-s,y).$
Thus, we get that
$\Lambda^{(3)}_{\alpha}(s,t,y)\lesssim(2-\alpha)\left[1\wedge\left(\|\sigma_{1}-\sigma_{2}\|(1+(t-s)^{-1/\alpha}|y|)\right)\right](t-s)^{-1}\varrho_{\alpha}(t-s,y).$
Combining the above calculations and observing (3.18), we finish the proof. ∎
@@
Let $\gamma_{1}\in[\alpha_{0},2]$ and $\gamma_{2}\in[\alpha_{0},2)$ with some $\alpha_{0}\in(1,2)$. There is a constant $c>0$ only depends on $d,\alpha_{0},\kappa_{1}$ such that for all $0\leq s<t<\infty$,
$\quad\int_{(t+s)/2}^{t}\int_{{\mathbb{R}}^{d}}|p^{(\gamma_{1}),\sigma}|(r,t,x-y)|\widetilde{\mathscr{D}}^{(\gamma_{2}),\sigma}_{r}p^{(\gamma_{2}),\sigma}|(s,r,y){\mathord{{\mathrm{d}}}}y{\mathord{{\mathrm{d}}}}r$
$\leq c(2-\gamma_{2})(1+|\ln(t-s)|)\left(1+|t-s|^{1-\frac{2}{\gamma_{2}}}\right)\sum_{i,j\in\{1,2\}}\varrho_{\gamma_{i},\gamma_{j}}(t-s,x),$
@@
By the second inequality of Lemma 3.5, one sees that
$\qquad|(\widetilde{{\mathscr{L}}}^{(\gamma_{2}),\sigma}_{r}-\widetilde{{\mathscr{L}}}^{(2),\sigma}_{r})p^{(\gamma_{2}),\sigma}(s,r,y)|$
@@
$\int_{{\mathbb{R}}^{d}}|p^{(\gamma_{1}),\sigma}|(r,t,x-y)|(\widetilde{{\mathscr{L}}}^{(\gamma_{2}),\sigma}_{r}-\widetilde{{\mathscr{L}}}^{(2),\sigma}_{r})p^{(\gamma_{2}),\sigma}|(s,r,y){\mathord{{\mathrm{d}}}}y$
@@
$\lesssim$ $~(2-\gamma_{2})(1+|\ln(t-s)|)(|t-s|^{-1}+|t-s|^{-\frac{2}{\gamma_{2}}})\sum_{i,j\in\{1,2\}}\varrho_{\gamma_{i},\gamma_{j}}(t-s,x).$
The desired estimate follows. ∎
$3.$ $\alpha$-continuity of heat kernels
@@
Assume that $\alpha\in[\alpha_{0},2]$ with some $\alpha_{0}\in(1,2)$. Then there is a constant $c=c(d,\alpha_{0},\kappa_{1})>0$ such that
@@
$\begin{split}|\nabla^{k}p^{(\alpha),\sigma}(s,s+t,x)-\nabla^{k}p^{(2),\sigma}(s,s+t,x)|\lesssim\varepsilon^{-1}(2-\alpha)t^{\frac{\alpha-2-\varepsilon}{\alpha}}\sum_{k_{1}=1}^{k}\sum_{\gamma_{1},\gamma_{2}\in\{\alpha,2\}}|x|^{2k_{1}-k}\varrho^{(2k_{1})}_{\gamma_{1},\gamma_{2}}(t,x).\end{split}$
@@
(i) Notice that the densities are even functions. From the defnition (3.4), one see that the following change of variable holds:
@@
$=\left(\int_{s}^{(t+s)/2}+\int^{t}_{(t+s)/2}\right){\langle}p_{r,t}^{(2),\sigma}(x-\cdot),(\widetilde{{\mathscr{L}}}_{r}^{(\alpha),\sigma}-\widetilde{{\mathscr{L}}}_{r}^{(2),\sigma})p_{s,r}^{(\alpha),\sigma}(\cdot){\rangle}{\mathord{{\mathrm{d}}}}r$
@@
$\quad+\int^{t}_{(t+s)/2}{\langle}p_{r,t}^{(2),\widetilde{\sigma}}(x-\cdot),(\widetilde{{\mathscr{L}}}_{r}^{(\alpha),\widetilde{\sigma}}-\widetilde{{\mathscr{L}}}_{r}^{(2),\widetilde{\sigma}})p_{s,r}^{(\alpha),\widetilde{\sigma}}(\cdot){\rangle}{\mathord{{\mathrm{d}}}}r$
with $\widetilde{\sigma}(r)=\sigma(r+(t-s)/2)$. Then (3.21) follows by Corollary 3.6.
@@ 4.2. Parametrix expansion of densities
$\mathfrak{B}_{s}^{(\tau,\xi)}f(x):=\left(b(x)-b(\theta^{(1)}_{\tau,s}(\xi)\right)\cdot\nabla f(x).$
@@
$\leq C^{n+1}\frac{{\Big(}\Gamma(\eta_{2}^{\varepsilon}){\Big)}^{n+1}}{\Gamma((n+1)\eta_{2}^{\varepsilon})}\sum_{\gamma_{1},\gamma_{2}\in\{\alpha,2\}}\phi^{(\eta_{1}^{\varepsilon}+n\eta_{2}^{\varepsilon})}_{\gamma_{1},\gamma_{2}}$.
Noticing that
$\phi^{(\eta_{1}^{\varepsilon}+n\eta_{2}^{\varepsilon})}_{\gamma_{1},\gamma_{2}}(s,t;x,y)\leq(1+T)^{\frac{2-\alpha_{0}}{\alpha_{0}}}\phi^{((n+1)\eta_{2}^{\varepsilon})}_{\gamma_{1},\gamma_{2}}(s,t;x,y),$
@@
$|q^{(\alpha)}(s,t;x,y)-q^{(2)}(s,t;x,y)|\leq c\varepsilon^{-1}(2-\alpha)\sum_{\gamma_{1},\gamma_{2}\in\{\alpha,2\}}\phi^{(\eta_{1}^{\varepsilon})}_{\gamma_{1},\gamma_{2}}(s,t;x,y).$
@@
${\boldsymbol{p}}_{b}^{(\alpha)}$ $={\boldsymbol{p}}_{0}^{(\alpha)}+{\boldsymbol{p}}_{0}^{(\alpha)}\otimes q^{(\alpha)}\lesssim{\Big(}\phi^{(1)}_{\alpha}+\phi^{(1)}_{\alpha}\otimes\phi^{(\eta_{0})}_{\alpha}{\Big)}(s,t;x,y)$
@@
$|({\boldsymbol{p}}_{0}^{(\alpha)}-{\boldsymbol{p}}_{0}^{(2)})\otimes q^{(2)}|$ $\lesssim\varepsilon^{-1}(2-\alpha)\sum_{\gamma_{1},\gamma_{2}\in\{\alpha,2\}}\left(\phi^{(\frac{2\alpha-2}{\alpha}-\frac{\varepsilon}{\alpha})}_{\gamma_{1},\gamma_{2}}\otimes\phi^{(\frac{\beta+1}{2})}_{2}\right)$
@@
$({{\mathscr{L}}}^{(\alpha)}+b\cdot\nabla)f(x)\lesssim\sup_{\alpha\in[\frac{\gamma+2}{2},2)}\int_{{\mathbb{R}}^{d}}(|z|^{2}{\boldsymbol{1}}_{|z|\leq 1}+|z|^{\gamma}{\boldsymbol{1}}_{|z|>1})\nu^{(\alpha)}({\mathord{{\mathrm{d}}}}z)+1-|x|^{r+\gamma}\leq 2cc_{0}-2c_{0}|x|^{r+\gamma}.$
@@
${\boldsymbol{d}}(\mu^{(\alpha)},\mu^{(2)})={\boldsymbol{d}}({P_{t}^{(\alpha)}}^{*}\mu^{(\alpha)},{P_{t}^{(2)}}^{*}\mu^{(2)})\leq{\boldsymbol{d}}({P_{t}^{(\alpha)}}^{*}\mu^{(\alpha)},{P_{t}^{(2)}}^{*}\mu^{(\alpha)})+{\boldsymbol{d}}({P_{t}^{(2)}}^{*}\mu^{(\alpha)},{P_{t}^{(2)}}^{*}\mu^{(2)}),$

### 26 Mar 2026 v2

@@ 4.2. $\alpha$-stability estimates of infinitesimal generators
$|\Lambda^{(2)}_{\alpha}(s,t,y)|\lesssim(2-\alpha)|t-s|^{-1}\varrho_{\alpha}(t-s,y).$
(iii) Under the condition (4.1), by Corollary 4.4, we have
$|\delta_{\sigma_{1}z}p^{(2),\sigma}_{s,t}(y)|\lesssim|t-s|^{\frac{\alpha-2}{\alpha}}{\Big(}\varrho_{\alpha}(t-s,y+\sigma_{1}z)+\varrho_{\alpha}(t-s,y){\Big)}.$
Thus,
$\Lambda^{(3)}_{\alpha}(s,t,y$ $)\lesssim(2-\alpha)|t-s|^{\frac{\alpha-2}{\alpha}}|t-s|^{-1}\varrho_{\alpha}(t-s,y)$
$\,=(2-\alpha)|t-s|^{-\frac{2}{\alpha}}\varrho_{\alpha}(t-s,y).$
Combining the above calculations and observing (4.17), we get the result for $\gamma=2$.
The proof is finished. ∎
Remark 4.7.
Following the proof of the lemma above, one sees that: Assume that $\sigma_{1},\sigma$ satisfy the condition (4.1). If $\alpha\in[\alpha_{0},2)$ with some $\alpha_{0}\in(0,2)$, then there is a constant $c=c(d,\kappa_{1},\alpha_{0})>0$ such that for $0\leq s<r<t<+\infty$,
$|\widetilde{\mathscr{D}}^{(\alpha),\sigma_{1}}_{r}p^{(\alpha),\sigma}|(s,t,y)$ $\leq c(2-\alpha)(1+|\ln(t-s)|)(|t-s|^{-1}+|t-s|^{-\frac{2}{\alpha}})\varrho_{\alpha}(t-s,y),$
where $\widetilde{\mathscr{D}}^{(\alpha),\sigma}_{r}$ is defined by (4.16).
@@
Fix $T>0$. Assume that $\sigma_{1},\sigma_{2},\sigma$ satisfy the condition (4.1). Let $\gamma_{1}\in[\alpha_{0},2]$ and $\gamma_{2}\in[\alpha_{0},2)$ with some $\alpha_{0}\in(0,2)$. There is a constant $c>0$ depending only on $d,T,\alpha_{0},\kappa_{1}$ such that for all $0\leq s<t\leq T$,
$\quad\int_{(t+s)/2}^{t}\int_{{\mathbb{R}}^{d}}|p^{(\gamma_{1}),\sigma_{1}}|(r,t,x-y)|\widetilde{\mathscr{D}}^{(\gamma_{2}),\sigma}_{r}p^{(\gamma),\sigma_{2}}|(s,r,y){\mathord{{\mathrm{d}}}}y{\mathord{{\mathrm{d}}}}r$
$\leq c(2-\gamma_{2})(1+|\ln(t-s)|)(|t-s|^{-1}+|t-s|^{-\frac{2}{\gamma_{2}}})\sum_{i,j\in\{1,2\}}\varrho_{\gamma_{i},\gamma_{j}}(t-s,x)$
where $\gamma=\{\gamma_{2},2\}$ and $\widetilde{\mathscr{D}}^{(\gamma_{2}),\sigma}_{r}$ is defined by (4.16).
@@
By Lemma 4.6, one sees that
$\qquad|(\widetilde{{\mathscr{L}}}^{(\gamma_{2}),\sigma}_{r}-\widetilde{{\mathscr{L}}}^{(2),\sigma}_{r})p^{(\gamma),\sigma_{2}}(s,r,y)|$
@@
$\int_{{\mathbb{R}}^{d}}|p^{(\gamma_{1}),\sigma_{1}}|(r,t,x-y)|(\widetilde{{\mathscr{L}}}^{(\gamma_{2}),\sigma}_{r}-\widetilde{{\mathscr{L}}}^{(2),\sigma}_{r})p^{(\gamma),\sigma_{2}}|(s,r,y){\mathord{{\mathrm{d}}}}y$
@@
$\lesssim$ $~(2-\gamma_{2})(1+|\ln(t-s)|)(|t-s|^{-1}+|t-s|^{-\frac{2}{\gamma_{2}}})\sum_{i,j\in\{1,2\}}\varrho_{\gamma_{i},\gamma_{j}}(t-s,x),$
where we used (3.23) in the last inequality. The desired estimate follows. ∎
Remark 4.9.
Following the proof of the corollary above, one sees that: Assume that $\sigma_{1},\sigma_{2},\sigma$ satisfy the condition (4.1). Let $\gamma_{1}\in[\alpha_{0},2]$ and $\gamma_{2}\in[\alpha_{0},2)$ with some $\alpha_{0}\in(0,2)$. There is a constant $c>0$ depending only on $d,\alpha_{0},\kappa_{1}$ such that for all $0\leq s<t<\infty$,
$\quad\int_{(t+s)/2}^{t}\int_{{\mathbb{R}}^{d}}|p^{(\gamma_{1}),\sigma_{1}}|(r,t,x-y)|\widetilde{\mathscr{D}}^{(\gamma_{2}),\sigma}_{r}p^{(\gamma_{2}),\sigma_{2}}|(s,r,y){\mathord{{\mathrm{d}}}}y{\mathord{{\mathrm{d}}}}r$
$\leq c(2-\gamma_{2})(1+|\ln(t-s)|)\left(|t-s|^{-1}+|t-s|^{-\frac{2}{\gamma_{2}}}\right)\sum_{i,j\in\{1,2\}}\varrho_{\gamma_{i},\gamma_{j}}(t-s,x)$
4.3. $\alpha$-continuity of heat kernels
@@
Assume that $\alpha\in[\alpha_{0},2]$ with some $\alpha_{0}\in(0,2)$. Then there is a constant $c=c(d,T,\alpha_{0},\kappa_{1})>0$ such that such that for any $0\leq s<s+t\leq T$,
@@
$\quad|\nabla^{k}p^{(\alpha),\sigma}(s,s+t,x)-\nabla^{k}p^{(2),\sigma}(s,s+t,x)|$
$\lesssim\varepsilon^{-1}(2-\alpha)t^{\frac{\alpha-2-\varepsilon}{\alpha}}\sum_{k_{1}=1}^{k}\sum_{\gamma_{1},\gamma_{2}\in\{\alpha,2\}}|x|^{2k_{1}-k}\varrho^{(2k_{1})}_{\gamma_{1},\gamma_{2}}(t,x).$
@@
(i) Notice that the densities are even functions. From the defnition (4), one see that the following change of variable holds:
@@
$\quad+\int_{s}^{(t+s)/2}{\langle}p_{s,r}^{(\alpha),\sigma}(\cdot),(\widetilde{{\mathscr{L}}}_{r}^{(\alpha),\sigma}-\widetilde{{\mathscr{L}}}_{r}^{(2),\sigma})p_{r,t}^{(2),\sigma}(x-\cdot){\rangle}{\mathord{{\mathrm{d}}}}r$
$:=I_{1}+I_{2}$
Moreover, notice that, by change of variables and the scaling property (4.3), we get that
$I_{2}$ $=\int_{(t+s)/2}^{t}{\langle}p_{s,t+s-r}^{(\alpha),\sigma}(\cdot),(\widetilde{{\mathscr{L}}}_{t+s-r}^{(\alpha),\sigma}-\widetilde{{\mathscr{L}}}_{t+s-r}^{(2),\sigma})p_{t+s-r,t}^{(2),\sigma}(x-\cdot){\rangle}{\mathord{{\mathrm{d}}}}r$
$=\int_{(t+s)/2}^{t}{\langle}p_{r,t}^{(\alpha),\sigma_{2}},(\widetilde{{\mathscr{L}}}_{r}^{(\alpha),\widetilde{\sigma}}-\widetilde{{\mathscr{L}}}_{r}^{(2),\widetilde{\sigma}})p_{s,r}^{(2),\sigma_{1}}(x-\cdot){\rangle}{\mathord{{\mathrm{d}}}}r,$
where
$\widetilde{\sigma}(u)=\sigma(t+s-u),\quad\sigma_{1}(u)=\sigma(t-r+u),\quad\sigma_{2}(u)=\sigma(s-r+u).$
Applying Corollary 4.8 to $I_{1},I_{2}$, (4.18) follows.
@@ 5.2. Parametrix expansion of densities
$\mathfrak{B}_{s}^{(\tau,\xi)}f(x):=\left(b(x)-b(\theta^{(1)}_{\tau,s}(\xi))\right)\cdot\nabla f(x).$
@@
$\leq C^{n+1}\frac{{\Big(}\Gamma(\eta_{2}^{\varepsilon}){\Big)}^{n+1}}{\Gamma((n+1)\eta_{2}^{\varepsilon})}\sum_{\gamma_{1},\gamma_{2}\in\{\alpha,2\}}\phi^{(\eta_{1}^{\varepsilon}+n\eta_{2}^{\varepsilon})}_{\gamma_{1},\gamma_{2}}$,
@@
$|q^{(\alpha)}-q^{(2)}|(s,t;x,y)\leq c\varepsilon^{-1}(2-\alpha)\sum_{\gamma_{1},\gamma_{2}\in\{\alpha,2\}}\phi^{(\eta_{1}^{\varepsilon})}_{\gamma_{1},\gamma_{2}}(s,t;x,y).$
@@
${\boldsymbol{p}}_{b}^{(\alpha)}$ $={\boldsymbol{p}}_{0}^{(\alpha)}+{\boldsymbol{p}}_{0}^{(\alpha)}\otimes q^{(\alpha)}\lesssim{\Big(}\phi^{(1)}_{\alpha}+\phi^{(1)}_{\alpha}\otimes\phi^{(\eta_{0})}_{\alpha}{\Big)}$
@@
$\quad|({\boldsymbol{p}}_{0}^{(\alpha)}-{\boldsymbol{p}}_{0}^{(2)})\otimes q^{(2)}|$
$\lesssim\varepsilon^{-1}(2-\alpha)\sum_{\gamma_{1},\gamma_{2}\in\{\alpha,2\}}\left(\phi^{(\frac{2\alpha-2}{\alpha}-\frac{\varepsilon}{\alpha})}_{\gamma_{1},\gamma_{2}}\otimes\phi^{(\frac{\beta+1}{2})}_{2}\right)$
@@
$({{\mathscr{L}}}^{(\alpha)}+b\cdot\nabla)f(x)$ $\lesssim\sup_{\alpha\in[\frac{\gamma+2}{2},2)}\int_{{\mathbb{R}}^{d}}(|z|^{2}{\boldsymbol{1}}_{|z|\leq 1}+|z|^{\gamma}{\boldsymbol{1}}_{|z|>1})\nu^{(\alpha)}({\mathord{{\mathrm{d}}}}z)+1-|x|^{r+\gamma}$
$\leq 2cc_{0}-2c_{0}|x|^{r+\gamma}.$
@@
${\boldsymbol{d}}(\mu^{(\alpha)},$ $\mu^{(2)})={\boldsymbol{d}}({P_{t}^{(\alpha)}}^{*}\mu^{(\alpha)},{P_{t}^{(2)}}^{*}\mu^{(2)})$
$\leq{\boldsymbol{d}}({P_{t}^{(\alpha)}}^{*}\mu^{(\alpha)},{P_{t}^{(2)}}^{*}\mu^{(\alpha)})+{\boldsymbol{d}}({P_{t}^{(2)}}^{*}\mu^{(\alpha)},{P_{t}^{(2)}}^{*}\mu^{(2)}),$